# Gravity and fluid dynamic correspondence on a null hypersurface: inconsistencies and advancement

### v1

@@ Appendix C Energy conservation relation in null hydrodynamics
The projection component $R_{ab}l^{a}l^{b}$ is related to the dynamical evolution of the expansion scalar corresponding to the null generators $\bm{{l}}$. This is essentially the null Raychaudhuri equation [25, 49], In the absence of matter, and expanding the L.H.S of Eq. (47), we have the NRE as,
$l^{a}\nabla_{a}\theta-\kappa\theta+\frac{1}{2}\theta^{2}+\sigma_{ab}\sigma^{ab}=0~{}.$ (C.1)
@@
Since the evolution of the expansion scalar is considered along the null generators of $\mathcal{H}$, we have in the adapted coordinate system, $l^{\mu}\overset{\mathcal{H}}{=}(1,V^{A})$. Thus expanding $l^{a}\nabla_{a}\theta=\partial_{t}\theta+V^{A}\partial_{A}\theta$, we have from (C.1),
$-\partial_{t}\theta-v^{A}\partial_{A}\theta+\kappa\theta-\theta^{2}=\sigma_{ab}\sigma^{ab}-\frac{1}{2}\theta^{2}~{}.$ (C.2)
@@
Now, we consider the specific case in which the induced metric on the submanifold $(\mathcal{B}^{(null)},\bm{q},\bm{~{}^{(2)}D})$ is stationary or independent of the time evolution parameter $t$, i.e. $\partial_{t}q_{AB}\overset{\mathcal{H}}{=}0$. Then from (24), we have that the second fundamental form takes the structure,
$\Theta_{AB}\overset{\mathcal{H}}{=}\frac{1}{2}\Big{(}~{}^{(2)}D_{A}V_{B}+~{}^{(2)}D_{B}V_{A}\Big{)}~{},$ (C.3)
@@
with the expansion scalar being given as $\theta\overset{\mathcal{H}}{=}~{}^{(2)}D_{A}V^{A}$. Inserting this in (C.2), we have,
$\partial_{t}(-\theta)+V^{A}\partial_{A}(-\theta)+(\kappa-\theta)~{}^{(2)}D_{A}V^{A}=\sigma_{ab}\sigma^{ab}-\frac{1}{2}(~{}^{(2)}D_{A}V^{A})^{2}~{}.$ (C.4)
Dividing (C.4) by a factor of $1/8\pi$ essentially leads us to (48).

### v2

@@ Appendix D Energy conservation relation in null hydrodynamics
The projection component $R_{ab}l^{a}l^{b}$ is related to the dynamical evolution of the expansion scalar corresponding to the null generators $\bm{{l}}$. This is essentially the null Raychaudhuri equation [25, 54], In the absence of matter, and expanding the L.H.S of Eq. (Eq. (), we have the NRE as,
$l^{a}\nabla_{a}\theta-\kappa\theta+\frac{1}{2}\theta^{2}+\sigma_{ab}\sigma^{ab}=0~.$ (D.1)
@@
Since the evolution of the expansion scalar is considered along the null generators of $\mathcal{H}$, we have in the adapted coordinate system, $l^{\mu}\overset{\mathcal{H}}{=}(1,V^{A})$. Thus expanding $l^{a}\nabla_{a}\theta=\partial_{t}\theta+V^{A}\partial_{A}\theta$, we have from (Eq. (),
$-\partial_{t}\theta-v^{A}\partial_{A}\theta+\kappa\theta-\theta^{2}=\sigma_{ab}\sigma^{ab}-\frac{1}{2}\theta^{2}~.$ (D.2)
@@
Now, we consider the specific case in which the induced metric on the submanifold $(\mathcal{B}^{\rm(null)},\bm{q},\bm{~{}^{(2)}D})$ is stationary or independent of the time evolution parameter $t$, i.e. $\partial_{t}q_{AB}\overset{\mathcal{H}}{=}0$. Then from (Eq. (), we have that the second fundamental form takes the structure,
$\theta_{AB}\overset{\mathcal{H}}{=}\frac{1}{2}\Big(~^{(2)}D_{A}V_{B}+~^{(2)}D_{B}V_{A}\Big)~,$ (D.3)
@@
with the expansion scalar being given as $\theta\overset{\mathcal{H}}{=}~^{(2)}D_{A}V^{A}$. Inserting this in (Eq. (), we have,
$\partial_{t}(-\theta)+V^{A}\partial_{A}(-\theta)+(\kappa-\theta)~^{(2)}D_{A}V^{A}=\sigma_{ab}\sigma^{ab}-\frac{1}{2}(~^{(2)}D_{A}V^{A})^{2}~.$ (D.4)
Dividing (Eq. () by a factor of $1/8\pi$ essentially leads us to (Eq. ().